OpenStudy (anonymous):

First one to answer get a prize. A "bit" is 1/8 of a dollar. How many bits are in a dollar.

OpenStudy (anonymous):

8 my inspection and the definition of a bit.

OpenStudy (anonymous):

8

OpenStudy (anonymous):

Can you tell me how to get to that answer?

OpenStudy (anonymous):

yeah.It is like this... 1*8,that simple.

OpenStudy (anonymous):

Divide 1 by 1/8.

OpenStudy (anonymous):

yeah but the dividing format is like this: 1*8 because of the reciprocal

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

no probs ^^

OpenStudy (anonymous):

\[\frac{1}{\frac{1}{8}}=1*\frac{8}{1}=8 \]